Molander Corporation is a distributor of a sun umbrella used at resort hotels. Data concerning the next month’s budget appear below:

Selling price per unit $ 23
Variable expense per unit $ 17
Fixed expense per month $ 4,980
Unit sales per month 980

Required:

1. What is the company’s margin of safety? (Do not round intermediate calculations.)

2. What is the company’s margin of safety as

1. Margin of safety (in dollars)
2. Margin of safety percentage %
Answer #1
Fixed expense per month 4980
Divide by Unit Contribution margin 6 =23-17
Break even sales units 830
Margin of safety units 150 =980-830
1
Margin of safety (in dollars) 3450 =150*23
2
Margin of safety percentage 15.31% =150/980